# Contextualize-then-Aggregate: Circuits for In-Context Learning in Gemma-2 2B

### v1

@@ Appendix A ICL on Linear Regression Benefits from Contextualization Step
A substantial amount of work has studied ICL in small transformers trained to solve simple synthetic tasks, particularly linear regression. However, the link between these mechanisms and the ICL circuits inside LLMs largely remain unknown. Theoretical constructions for ICL on linear regression based on gradient descent (GD) and similar algorithms often involve just multiple rounds of parallel aggregation from each fewshot; in particular, a single round of GD can be implemented as a single aggregation step, and is the optimal ICL strategy for a one-layer (linear) transformer (Mahankali et al., 2023). However, there is theoretical reason to believe that a contextualization step is beneficial in preprocessing the representations of individual fewshots before aggregation. First, multilayer linear transformers implement the GD++ algorithm (Von Oswald et al., 2023), which iteratively contextualizes $x_{i}$’s to approximate the input data covariance (Von Oswald et al., 2023), and which allows faster convergence than GD (Vladymyrov et al., 2024); such preconditioning also appears in constructions for ICL on non-isotropic data (Ahn et al., 2023). Second, Chen et al. (2024) showed that preprocessing of individual fewshots is beneficial in sparse linear regression. While these works show that contextualization helps outperform gradient descent, they do not establish that the step is needed for optimal performance, as a priori there might be undiscovered algorithms implementable by transformers outperforming GD with only the aggregation step.
We empirically investigated which attention edges are needed when transformers learn linear regression in context. Following Garg et al. (2022), we trained transformers to perform linear regression, i.e., each prompt has a latent task $v\in\mathbb{R}^{d}$ and each fewshot consists of a vector $x\in\mathbb{R}^{d}$ and the output $y=v^{T}x\in\mathbb{R}$. We chose $d=20$.
@@
    Parallel: $x_{i}\rightarrow y_{i}$, $y_{i}\rightarrow x_{N+1}$
@@
Results are in Figure 6. The Parallel subcircuit performs much better than random baseline (MSE loss of 20), but suffers substantially for shallow models. $x_{i}\rightarrow x_{j}$ edges improve, but cannot close the gap to the full model. $y_{i}\rightarrow y_{j}$ edges nearly or fully close the gap. Overall, while parallel aggregation plays an important role, a contextualization step is needed to close the gap to the full transformer.
@@
Figure 6: Loss (MSE) on ICL on linear functions when trained with attention links ablated during training, with transformers with $l=4,6,8,12$ layers. The Parallel circuit performs much better than a random baseline (loss 20.0), but underperforms the full model by a large margin in shallow transformers. Adding links within $x$’s or $y$’s improves performance, particularly the latter. Adding both performs at par with the full model, in line with our findings on real-world tasks in Gemma-2.
@@ Appendix C Datasets
All datasets were randomly divided into train and test parts. To construct an ICL prompt, we select examples from the train split of the dataset without replacement to use as fewshot examples, and one example from the test split to use as final query and target. We used a test size of 100 examples in each dataset. [^9] The high computational cost of the full set of patching experiments (Figure 8) prevented us from running larger test sets. Importantly, our key conclusions are drawn only from patching effects with statistically significant changes in accuracy (Figure. 8. Train sizes of datasets range from 39 to 4000.
@@
Appendix D Overall Set of Edges
@@
Note that our visualizations of circuits only include positions from $x_{1}$ to $t_{N+1}$, as positions before or after are not connected to the $N$ fewshot examples by any edges; recall that our focus is on the process by which information is assembled from the fewshot examples.
Appendix E Details for Gradient-Based Circuit Finding Algorithm
@@
[…]
@@
E.1 Step 1: Head pruning
@@
[…]
@@
E.2 Step 2: Edge pruning
@@
We distinguish between 16 types of tokens: 3 types for queries in the fewshot examples, 3 types for targets in fewshot examples, 3 types for separators between queries and targets in fewshot examples, 3 types for separators between fewshot examples, a type for query of the whole input, a type for target that the model needs to generate, a type for the last separator before target, a type for bos/eos/pad tokens. In Figure 1 we use color coding to differentiate token types, though some are merged into single categories for visual simplicity. We view each activation site as 16 nodes that correspond to different token types.
@@
We stop pruning edges based on a threshold on increased loss ($1.5*loss\_before\_pruning$) and decreased accuracy ($0.9*accuracy\_before\_pruning$). As above, thresholds were chosen in preliminary experiments and not extensively tuned; these could be refined in future work.
Cap CC PP Copy PS
@@
Table 1: Accuracy of Gemma2-2B on circuits when adding edges between different xs and ys combinations to the Parallel-only circuit. Our aim is to find a minimal circuit sufficient for recovering most ($\approx 90\%$) of the full model’s accuracy. Adding (2) most consistently improves accuracy compared to the Parallel-only circuit, with a large increase on Capitalization, where the other types of edges achieve no such increase. Adding (1) or (4) both substantially improve accuracy over the Parallel-only circuit in Country-Capital. We note that adding (1) results in a smaller number of possible paths than (4), which simplifies interpretation of information flow; hence, focused on (1). In other words, to keep the circuits tractable and understandable, we do not aim to guarantee completeness, but aim to provide minimal and sufficient circuits. Overall, we focused on (1) and (2) for determining position-level circuits. As we show, our resulting position-level circuits are indeed sufficient for recovering $\geq 90\%$ of the full model’s accuracy (Figure 2).
Full remove-seps Parallel Parallel Parallel
@@
Parallel refers to the circuit consisting of only $x_{i}\rightarrow y_{i}$, $y_{i}\rightarrow t_{N+1}$, $x_{N+1}\rightarrow t_{N+1}$ edges.
$y_{i}\rightarrow y_{j}$ refers to the edges $y_{i}\rightarrow y_{j}$ ($11\leq i<j\leq N-$ 1).
$y_{i}\rightarrow y_{i+1}$ refers to the edges $y_{i}\rightarrow y_{i+1}$ ($11\leq i\leq N-$ 1).
@@
Table 5: Accuracy results of circuits on Gemma2-9B. Removing edges between the separators hurts even less than in the 2B model (compare Table 2 for that). As in the 2B model, accuracy of the Parallel-only circuit is low on some tasks, in particular Capitalization. Including contextualization recovers almost the entire performance of the full model. In bold are the circuits we had selected based on the 2B model. Compare Table 2 for the 2B model, and Table 6 for the 27B model.
PP CC Copy Cap PS
@@
Appendix F Error Analysis
@@
Capitalization (Parallel Subcircuit)
@@
Figure 14: Error patterns: 10 shots (top) and 3 shots (bottom), $y_{i}\rightarrow t_{N+1}$(Capitalization). Perturbing the functional behavior but leaving the output space intact leads to incorrect responses staying in the correct output space (capitalized tokens; dark blue). Disrupting the output space leads to production of other tokens. See Appendix F for further information.
Copying (Parallel Subcircuit)
@@
Figure 15: Error patterns for ablations in the (aggregation-only) circuit for Copying. Breaking the functional relationship between $x$ and $y$ often leads to reproduction of other tokens from the corrupted prompt (light green), or the production of other tokens (yellow). See Appendix F for further information.
Country-Capital (Parallel Subcircuit)
@@
Figure 16: Error patterns for ablations in the aggregation edges for Country-Capital. See Appendix F for further information.
Person-Sport (Parallel Subcircuit)
@@
Figure 17: Error patterns in the Person-Sport task. The downstream effect of the aggregation edges suffers only weakly unless $x$ or $y$ leave their respective spaces. If $x$ leaves $\mathcal{X}$, predictions still are often correct or from the correct output space, showing substantial robustness in task inference. If $y$ leaves $\mathcal{Y}$, the model copies a token from the prompt or provides some other token. See Appendix F for further information.
Present-Past (Parallel Subcircuit)
@@
Figure 18: Present-Past task. Aggregation subcircuit. Disrupting the functional relation between $x$ and $y$, even if leaving the input- and output-spaces unchanged, leads to a large amount of reproduction from the correct output space (past-tense verbs), copying from the prompt, or other errors. See Appendix F for further information.
@@
Figure 19: Error patterns for ablating $y_{i}\rightarrow y_{j}$ edges (Capitalization). The $y$-axis denotes the number of datapoints in each class. Ablating information about the output space leads to a large number of copy-type mistakes. See Appendix F for further information.
@@
Figure 20: Error patterns for ablations in the $x_{i}\rightarrow x_{i+1}$ edges for Country-Capital. The $y$-axis denotes the number of datapoints in each class. Corrupting the input-space information in these edges leads to a substantial fraction of copying responses, i.e., the functional input-output behavior breaks down, in particular at 3 shots. See Appendix F for further information.
@@
Figure 21: Error patterns for ablations in the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital. The $y$-axis denotes the number of datapoints in each class. Corrupting the output space information produces a (limited but statistically significant) increase in copying responses. See Appendix F for further information.
Appendix G More Details on Corrupted Prompts
@@ 7. Ablating examples, while preserving the functional relationship:
Figure 26: Top-10 function vector heads on each task, and their roles in the activation-level circuits (3-shot prompts, $N=3$). Edges from the form $y_{i}\rightarrow t_{N+1}$ are highlighted in boldface. Most are involved in at least one such edge, showing that function vector heads are causally involved in the aggregation of task information from fewshot examples. Many edges also causally participate in $t_{N+1}\rightarrow t_{N+1}$ edges, suggesting processing or forwarding of task information. Heads also sometimes participare in edges not going to $t_{N+1}$ (those do not enter the function vector score calculation); these are also shown here. For the two remaining tasks, see next figure.
@@
Table 8: Determining through which downstream paths $x_{i}\rightarrow x_{j}$ edges provides information: A priori, given the position-level circuit for the Country-Capital task (Figure 24), $x_{i}\rightarrow x_{j}$ edges might provide information affecting the final prediction both via $x_{i}\rightarrow x_{i+1}\rightarrow y_{i+1}\rightarrow t_{N+1}$ and via $x_{N}\rightarrow x_{N+1}\rightarrow t_{N+1}$. Here, we patch with prior $x_{i}$’s, for both types of donwstream edges ($x_{N+1}\rightarrow t_{N+1}$ and $y_{i}\rightarrow t_{N+1}$), on the Country-Capital task. We consider both patching with random words (eliminating information both about $x_{i}$ and $\mathcal{X}$), and patching with other words in the input space (eliminating information about $x_{i}$ but not $\mathcal{X}$). Given the position-level circuit for the Country-Capital task (Figure 24), information about the manipulated words can flow into these edges only via $x_{i}\rightarrow x_{i+1}$ connections. Accuracy drops substantively compared to the full circuit (0.79 at 3 shots, 0.8 at 10 shots) when eliminating information about $\mathcal{X}$, both when applying this patch to $x_{N+1}\rightarrow t_{N+1}$ or to $y_{i}\rightarrow t_{N+1}$. This shows that $x_{i}\rightarrow x_{i+1}$ edges contextualize both the individual fewshot examples and the query $x_{N+1}$ with information about the input space $\mathcal{X}$.
Full Parallel remove-seps Parallel
@@
Table 9: Accuracy of the Present-Past task in the presence of different numbers of ambiguous fewshot examples (Section 3.3). In the presence of ambiguity, the full model (Full) continues to perform well even when 7 out of 10 examples are ambiguous between the Present-Past and Copying tasks. In contrast, the accuracy of the Parallel circuit drops to 56% as the number of ambiguous examples increases from 0 to 7. Adding contextualization between $x_{i}$’s an $y_{i}$’s helps, but does not close the gap to the full model; indeed, even ablating edges involving separators preceding $t_{N+1}$ hurts (remove-seps), in contrast to the standard test sets.

### v3

@@ Appendix A ICL on Linear Regression Benefits from Contextualization Step
A substantial amount of work has studied ICL in small transformers trained to solve simple synthetic tasks, particularly linear regression. However, the link between these mechanisms and the ICL circuits inside LLMs largely remain unknown. Theoretical constructions for ICL on linear regression based on gradient descent (GD) and similar algorithms often involve just multiple rounds of parallel aggregation from each few-shot; in particular, a single round of GD can be implemented as a single aggregation step, and is the optimal ICL strategy for a one-layer (linear) transformer (Mahankali et al., 2023). However, there is theoretical reason to believe that a contextualization step is beneficial in preprocessing the representations of individual few-shots before aggregation. First, multilayer linear transformers implement the GD++ algorithm (Von Oswald et al., 2023), which iteratively contextualizes $x_{i}$’s to approximate the input data covariance (Von Oswald et al., 2023), and which allows faster convergence than GD (Vladymyrov et al., 2024); such preconditioning also appears in constructions for ICL on non-isotropic data (Ahn et al., 2023). Second, Chen et al. (2024) showed that preprocessing of individual few-shots is beneficial in sparse linear regression. While these works show that contextualization helps outperform gradient descent, they do not establish that the step is needed for optimal performance, as a priori there might be undiscovered algorithms implementable by transformers outperforming GD with only the aggregation step.
We empirically investigated which attention edges are needed when transformers learn linear regression in context. Following Garg et al. (2022), we trained transformers to perform linear regression, i.e., each prompt has a latent task $v\in\mathbb{R}^{d}$ and each few-shot consists of a vector $x\in\mathbb{R}^{d}$ and the output $y=v^{T}x\in\mathbb{R}$. We chose $d=20$.
@@
    Aggregation: $x_{i}\rightarrow y_{i}$, $y_{i}\rightarrow x_{N+1}$
@@
Results are in Figure 6. The Aggregation subcircuit performs much better than random baseline (MSE loss of 20), but suffers substantially for shallow models. $x_{i}\rightarrow x_{j}$ edges improve, but cannot close the gap to the full model. $y_{i}\rightarrow y_{j}$ edges nearly or fully close the gap. Overall, while parallel aggregation plays an important role, a contextualization step is needed to close the gap to the full transformer.
@@
Figure 6: Loss (MSE) on ICL on linear functions when trained with attention links ablated during training, with transformers with $l=4,6,8,12$ layers. The Aggregation circuit performs much better than a random baseline (loss 20.0), but underperforms the full model by a large margin in shallow transformers. Adding links within $x$’s or $y$’s improves performance, particularly the latter. Adding both performs at par with the full model, in line with our findings on real-world tasks in Gemma-2.
@@ Appendix C Datasets
All datasets were randomly divided into train and test parts. To construct an ICL prompt, we select examples from the train split of the dataset without replacement to use as few-shot examples, and one example from the test split to use as final query and target. We used a test size of 100 examples in each dataset. [^12] The high computational cost of the full set of patching experiments (Figure 8) prevented us from running larger test sets. Importantly, our key conclusions are drawn only from patching effects with statistically significant changes in accuracy (Figure. 8. Train sizes of datasets range from 39 to 4000.
@@
Appendix D Patching methodology
We use path patching (Wang et al., 2022; Hanna et al., 2023) to identify position-level circuits and analyze information flow along edges. This method involves: (1) representing the model as a computation graph with token embeddings as inputs, loss value as output, and specific activations as nodes; (2) ablating graph components to isolate those responsible for specific behaviors. This section details our computation graph definition (nodes, edges), loss function and ablation implementation.
Computation graph
Nodes consist of activations at the input/output of each attention head across positions. Each node is represented as a tuple (input/output, layer, head, position), where "position" indicates a token’s semantic role (e.g., $x_{1}$, $y_{3}$, $t_{2}$, $n_{4}$). Note that $x_{i}$ and $y_{i}$ positions typically span multiple tokens but are treated as single nodes. Each node is thus a tensor of size (number of tokens in position, hidden dimension).
[…]
instead of
[…]
where subscript $C$ denotes counterfactual activations.
When discovering position-level circuits, we abstract all edges between the same pair of positions into a single edge. Specifically, we treat all edges of the form ((input, $L$, $H$, pos=AA) $\rightarrow$ (output, $L$, $H$, pos=$B$) for all $L$, $H$) as a single edge: (input, pos=$A$) $\rightarrow$ (output, pos=$B$). Thus, removing such an edge corresponds to removing all edges between positions $A$ and $B$ across all attention heads and layers.
We use cross-entropy loss over all tokens in the output position ($y_{N+1}$). Performance drop is measured as accuracy decrease between ablated and normal model runs.
Ablation Procedure
Edge ablation requires two forward passes (Fig. 28):
  1. 1.
    Run counterfactual input $I_{C}$, saving all input node activations (input, L, H, pos)C
  2. 2.
    Run correct input $I$. For each attention head, compute outputs position-wise, which requires as many passed through attention block as there are positions. When ablating edge $A\rightarrow B$ in head $H$ (layer $L$), for position $B$: substitute activation at $A$ with (input, L, H, A)C saved from Step 1. Concatenate position outputs into final tensor.
With this methodology we can identify circuits - subgraphs of the model’s computation graph that can perform the task with high accuracy when the edges outside the circuit are ablated.
Our approach to circuit discovery is closely related to the recent work of Haklay et al. (2025), with several key differences:
  • The dataset schema in our case is defined by the specifics of the task: what we refer to as x’s, y’s, t’s, and n’s corresponds to what Haklay et al. (2025) call a “span.”
  • We define an edge from position $A$ to position $B$ as patching both the V and K activations at position $A$ when queried by position $B$, whereas Haklay et al. (2025) distinguish between edges originating from Q, K, and V activations.
  • We do not take edges from a position to itself into account and thus do not have separate nodes for MLPs, embeddings and logits.
  • To identify position-level circuits, we abstract all edges from position $A$ to position $B$ across all layers into a single edge. We only distinguish between layers when identifying activation-level circuits (Section 3.1).
For analyzing information routing through specific edge sets (Section 3.2), we partition edges into three categories: (1) Outside circuit, (2) Inside circuit but not under investigation, (3) Inside circuit and under investigation. The extended procedure adds a third step:
  1. 1.
    Run counterfactual input $I_{C}$, saving (input, L, H, pos)C
  2. 2.
    Run semi-counterfactual input $I_{SC}$, ablating edges outside circuit (Set 1). Save input node activations for heads corresponding to Set 3 edges: (input, L, H, pos)SC
  3. 3.
    Run correct input $I$ while ablating:
    • Set 1 edges using (input, L, H, pos)C
    • Set 3 edges using (input, L, H, pos)SC
This isolates performance changes caused by removing only Set 3 information not present in the semi-counterfactual. Counterfactual inputs (e.g., replacing "Berlin" with "table") remove more information than semi-counterfactual (e.g., replacing "Berlin" with "Paris"), which preserve broader categorical attributes. This reveals which information aspects along specific edges impact downstream predictions.
Handling Multiple Tokens per Position
Our approach requires identical token counts per position in counterfactual and correct prompts, enforced during construction. When ablating edges from position $A$, we replace its entire activation tensor with the counterfactual version.
For projecting position-level circuits to activation-level circuits, we assign each edge a single importance score by summing scores across all constituent token-level edges. Position-level circuits require no such aggregation. See Appendix F for activation-level circuit details.
Appendix E Overall Set of Edges
@@
Note that our visualizations of circuits only include positions from $x_{1}$ to $t_{N+1}$, as positions before or after are not connected to the $N$ few-shot examples by any edges; recall that our focus is on the process by which information is assembled from the few-shot examples.
Appendix F Details for Gradient-Based Circuit Finding Algorithm
@@
    (Head Pruning) Starting from the position-level circuits, we view the model as a computational graph where nodes are the outputs of attention heads indexed by $\langle\text{position},\text{layer},\text{head}\rangle$, and prune heads.
@@
F.1 Step 1: Head pruning
@@
We first view the model as a computation graph where nodes are indexed by $\langle\text{position},\text{layer},\text{head}\rangle$ ($\text{position}\in\{x_{1},y_{1},\dots,x_{N},y_{N},x_{N+1},t_{N+1},y_{N+1}\},\text{layer}\in\{1,\dots,n_{layers}\},\text{head}\in\{1,\dots,n_{heads}\}$). Notably, our circuits also include $y_{N+1}$ even though in principle the last separator is where the answer is predicted, which is in order to account for multi-token outputs. The loss for gradient-based pruning is the entire multi-token CE on the answer.
@@
F.2 Step 2: Edge pruning
@@
We distinguish between 16 types of tokens: 3 types for queries in the few-shot examples, 3 types for targets in few-shot examples, 3 types for separators between queries and targets in few-shot examples, 3 types for separators between few-shot examples, a type for query of the whole input, a type for target that the model needs to generate, a type for the last separator before target, a type for bos/eos/pad tokens. In Figure 1 we use color coding to differentiate token types, though some are merged into single categories for visual simplicity. We view each activation site as 16 nodes that correspond to different token types.
@@
We stop pruning edges based on a threshold on increased loss (1.5loss_before_pruning1.5*loss\_before\_pruning) and decreased accuracy (0.9accuracy_before_pruning0.9*accuracy\_before\_pruning). As above, thresholds were chosen in preliminary experiments and not extensively tuned; these could be refined in future work.
Cap $C$ $P$ Copy PS
@@
Table 1: Accuracy of Gemma2-2B on circuits when adding edges between different xs and ys combinations to the Aggregation-only circuit. Our aim is to find a minimal circuit sufficient for recovering most ($\approx 90\%$) of the full model’s accuracy. Adding (2) most consistently improves accuracy compared to the Aggregation-only circuit, with a large increase on Capitalization, where the other types of edges achieve no such increase. Adding (1) or (4) both substantially improve accuracy over the Aggregation-only circuit in Country-Capital. We note that adding (1) results in a smaller number of possible paths than (4), which simplifies interpretation of information flow; hence, focused on (1). In other words, to keep the circuits tractable and understandable, we do not aim to guarantee completeness, but aim to provide minimal and sufficient circuits. Overall, we focused on (1) and (2) for determining position-level circuits. As we show, our resulting position-level circuits are indeed sufficient for recovering $\geq 90\%$ of the full model’s accuracy (Figure 2).
Full remove-seps Aggregate Aggregate Aggregate
@@
Aggregate refers to the circuit consisting of only $x_{i}\rightarrow y_{i}$, $y_{i}\rightarrow t_{N+1}$, $x_{N+1}\rightarrow t_{N+1}$ edges.
$y_{i}\rightarrow y_{j}$ refers to the edges $y_{i}\rightarrow y_{j}$ ($1\leq i<j\leq N-1$).
$y_{i}\rightarrow y_{i+1}$ refers to the edges $y_{i}\rightarrow y_{i+1}$ ($1\leq i\leq N-1$).
@@
Table 5: Accuracy results of circuits on Gemma2-9B. Removing edges between the separators hurts even less than in the 2B model (compare Table 2 for that). As in the 2B model, accuracy of the Aggregation-only circuit is low on some tasks, in particular Capitalization. Including contextualization recovers almost the entire performance of the full model. In bold are the circuits we had selected based on the 2B model. Compare Table 2 for the 2B model, and Table 6 for the 27B model.
$P$ $C$ Copy Cap PS
@@
Appendix G Error Analysis
@@
Capitalization (Aggregation Subcircuit)
@@
Figure 14: Error patterns: 10 shots (top) and 3 shots (bottom), $y_{i}\rightarrow t_{N+1}$(Capitalization). Perturbing the functional behavior but leaving the output space intact leads to incorrect responses staying in the correct output space (capitalized tokens; dark blue). Disrupting the output space leads to production of other tokens. See Appendix G for further information.
Copying (Aggregation Subcircuit)
@@
Figure 15: Error patterns for ablations in the (aggregation-only) circuit for Copying. Breaking the functional relationship between $x$ and $y$ often leads to reproduction of other tokens from the corrupted prompt (light green), or the production of other tokens (yellow). See Appendix G for further information.
Country-Capital (Aggregation Subcircuit)
@@
Figure 16: Error patterns for ablations in the aggregation edges for Country-Capital. See Appendix G for further information.
Person-Sport (Aggregation Subcircuit)
@@
Figure 17: Error patterns in the Person-Sport task. The downstream effect of the aggregation edges suffers only weakly unless $x$ or $y$ leave their respective spaces. If $x$ leaves $\mathcal{X}$, predictions still are often correct or from the correct output space, showing substantial robustness in task inference. If $y$ leaves $\mathcal{Y}$, the model copies a token from the prompt or provides some other token. See Appendix G for further information.
Present-Past (Aggregation Subcircuit)
@@
Figure 18: Present-Past task. Aggregation subcircuit. Disrupting the functional relation between $x$ and $y$, even if leaving the input- and output-spaces unchanged, leads to a large amount of reproduction from the correct output space (past-tense verbs), copying from the prompt, or other errors. See Appendix G for further information.
@@
Figure 19: Error patterns for ablating $y_{i}\rightarrow y_{j}$ edges (Capitalization). The $y$-axis denotes the number of datapoints in each class. Ablating information about the output space leads to a large number of copy-type mistakes. See Appendix G for further information.
@@
Figure 20: Error patterns for ablations in the $x_{i}\rightarrow x_{i+1}$ edges for Country-Capital. The $y$-axis denotes the number of datapoints in each class. Corrupting the input-space information in Neighboring-Xs edges leads to a substantial fraction of copying responses, i.e., the functional input-output behavior breaks down, in particular at 3 shots. See Appendix G for further information.
@@
Figure 21: Error patterns for ablations in the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital. The $y$-axis denotes the number of datapoints in each class. Corrupting the output space information produces a (limited but statistically significant) increase in copying responses. See Appendix G for further information.
Appendix H More Details on Corrupted Prompts
@@ 7. Ablating examples, while preserving the functional relationship:
Figure 26: Top-10 function vector heads on each task, and their roles in the activation-level circuits (3-shot prompts, $N=3$). Edges from the form $y_{i}\rightarrow t_{N+1}$ are highlighted in boldface. Most are involved in at least one such edge, showing that function vector heads are causally involved in the aggregation of task information from few-shot examples. Many edges also causally participate in $t_{N+1}\rightarrow t_{N+1}$ edges, suggesting processing or forwarding of task information. Heads also sometimes participare in edges not going to $t_{N+1}$ (those do not enter the function vector score calculation); these are also shown here. For the two remaining tasks, see next figure.
@@
Table 8: Determining through which downstream paths $x_{i}\rightarrow x_{j}$ edges provides information: A priori, given the position-level circuit for the Country-Capital task (Figure 24), $x_{i}\rightarrow x_{j}$ edges might provide information affecting the final prediction both via $x_{i}\rightarrow x_{i+1}\rightarrow y_{i+1}\rightarrow t_{N+1}$ and via $x_{N}\rightarrow x_{N+1}\rightarrow t_{N+1}$. Here, we patch with prior $x_{i}$’s, for both types of donwstream edges ($x_{N+1}\rightarrow t_{N+1}$ and $y_{i}\rightarrow t_{N+1}$), on the Country-Capital task. We consider both patching with random words (eliminating information both about $x_{i}$ and $\mathcal{X}$), and patching with other words in the input space (eliminating information about $x_{i}$ but not $\mathcal{X}$). Given the position-level circuit for the Country-Capital task (Figure 24), information about the manipulated words can flow into these edges only via $x_{i}\rightarrow x_{i+1}$ connections. Accuracy drops substantively compared to the full circuit (0.79 at 3 shots, 0.8 at 10 shots) when eliminating information about $\mathcal{X}$, both when applying this patch to $x_{N+1}\rightarrow t_{N+1}$ or to $y_{i}\rightarrow t_{N+1}$. This shows that $x_{i}\rightarrow x_{i+1}$ edges contextualize both the individual few-shot examples and the query $x_{N+1}$ with information about the input space $\mathcal{X}$.
Full Aggregation remove-seps Aggregation
@@
Table 9: Accuracy of the Present-Past task in the presence of different numbers of ambiguous few-shot examples (Section 3.3). In the presence of ambiguity, the full model (Full) continues to perform well even when 7 out of 10 examples are ambiguous between the Present-Past and Copying tasks. In contrast, the accuracy of the Aggregation circuit drops to 56% as the number of ambiguous examples increases from 0 to 7. Adding contextualization between $x_{i}$’s an $y_{i}$’s helps, but does not close the gap to the full model; indeed, even ablating edges involving separators preceding $t_{N+1}$ hurts (remove-seps), in contrast to the standard test sets.